OpenStudy (anonymous):

The point (5, 7) lies on the line -2x + 8y = 46. Determine if this point is a solution of the system -2x + 8y = 46 x – 2y = -9

OpenStudy (anonymous):

To check if a point is a solution, we can substitute the point to LHS of the equation and check if the result equals to RHS. For (5,7) and x-2y=-9 LHS=5-2(7)=-9 RHS=-9 LHS=RHS Therefore, (5,7) lies on x-2y=-9 Hence, it is a solution of the system.

OpenStudy (anonymous):

thanks for helping me figure this out cocotsoi but what do you mean by RHS? and LHS?

OpenStudy (anonymous):

RHS= right hand side LHS= left hand side

OpenStudy (anonymous):

okay I see....your super!!

OpenStudy (anonymous):

Thank you :D